(1 point) A rumor spreads through a school. Let y(t) be the fraction of the population that has heard the rumor at time t andAnswer asappp
Answer #1

SOLUTION ODE YIH) t-y dy ky (1-4) It to At 8 AM 108 10.801 1000x100 y (0) = 0.108 kdt ty y (1-4) INTEGRATE Skat Sviny Inlog/C, e 760)= ekt 1+C, ekt Y(0) = oolog ce → 0.108 ltceo с. 1+ -) 0.108 + 01108 G = G =) Ci- 0.1086 = 0,108 Ci (1-0.108) = 0.108